Math 142 (Sections 1&2)  University of South Carolina
D. Meade  Fall 1995

Homework Problems for Chapter 7 Transcendental Functions

Section Page Problems
7.1  333 # 3, 6, 8, 9, 13, 15, 19, 22, 23, 24  (CR # 1, 2, 4)
7.2  339 # 1, 4, 8, 9, 17, 26  (CR # 1, 3, 4)
7.3  346 # 3, 8, 11, 18, 23, 28, 29, 30, 36, 37  (CR # 1, 3, 4)
7.4  352 # 1, 3, 4, 7, 17, 18, 24, 25, 27, 31  (CR # 1, 4)
7.5  359 # 1, 2, 5, 6, 8, 27ac  (CR # 2, 3, 4)
7.6  366 # 1, 3, 5, 8, 11, 25, 26, 31  (CR # 1, 3)
7.7  372 # 1, 2, 4, 6, 7, 13, 18, 23, 25, 27, 28, 29, 31, 34  (CR # 3)
7.9  381 # 1–5, 10, 12–15, 18, 20–25, 27–33, 36, 38, 41  (CR # 1–20)

\underline{Notes}

• These problems will reinforce the main points presented in the text, and in the lectures. You should be sure you know how to work all of them. Note that references to examples in the text are provided for many problems. Don’t be afraid to use these hints.

• Some problems will require you to use knowledge learned in Math 141. If you don’t remember the appropriate technique, look it up!

• If you need more practice, there are many other problems in the text that are similar to the ones I have assigned.

• Section 7.9 is the chapter review. These problems will be excellent practice when preparing for the exam.

• If you don’t know how to do a problem, ask about it in recitation!